# ISOPO: Proximal policy gradients without pi-old

### v1

@@ 1.2 Refresher on the natural policy gradient
2 The ISOPO gradient update
@@ 2.0.3 Alternative sequence re-scalings
1"""alternativere-scalings
2Examplesdiscussed:
3p=-1,q=0,r=0(boundindividualKLcontributions)
4p=0,q=0,r=-2(scalarmultipleclosesttoNPG)
5"""
6def rescaling(grad, F_norm, p=-1, q=0, r=0):
@@
10
11# regularize
12def reg(val, regularization = 0.0):
13 # use e.g. EMA of trailing minibatch means for the regularization
@@
15
16# regularize the square of the value
17def reg2(val, reg = 0.0):
@@
2.1 Results
I implemented ISOPO in VeRL [verl]. The experiments shown are for gsm8k with the Qwen-3 0.6B model. All runs use the group-relative advantage estimator, and all runs are without KL penalty. The two baselines shown are GRPO and REINFORCE. Here, GRPO refers to using PPO clipping of the importance sampling ratio, and REINFORCE refers to the same algorithm with no clipping. Like REINFORCE, ISOPO is run without PPO clipping. Figure 4 shows a comparison between best-of-3 validation curves for REINFORCE (green), GRPO (blue), and ISOPO (red) with q=1q=-1 and r=2r=-2.
Figure 5 shows KL divergence from the initial policy and validation score at step 50 for ISOPO with p=1p=-1 (normalization of the Fisher norm estimate, shown in red) vs q=1q=-1 (Sequence-wise euclidean normalization, shown in yellow). While both normalizations increase the validation score, only the Fisher normalization has the effect of decreasing the KL drift. The normalizations do not use regularization.
Figure 5: KL-drift and validation score for non-interacting ISOPO vs GRPO and REINFORCE. Multiple runs are shown for each algorithm and the convex hulls are shaded for visual clarity. ISOPO is shown with two different settings: with Fisher normalization p=1p=-1, $q=r=0$ (red) and with sequence-wise Euclidian normalization q=1,p=r=0q=-1,p=r=0 (yellow). The p=1p=-1 setting decreases the KL drift whereas $q=1$ does not.
@@ 3 Interacting variant
4 Code
The code for ISOPO will be available at https://github.com/nilin/isopo.

### v2

@@ 1.2 Refresher on the natural policy gradient
2 The non-interacting ISOPO gradient update
@@ 2.0.3 Alternative sequence re-scalings
1# alternative re-scalings
2def rescaling(grad, F_norm, p=-1, q=0, r=0):
@@
6
7# regularize
8def reg(val, regularization = 0.0):
9 # use e.g. EMA of trailing minibatch means for the expectation
@@
11
12# regularize the square of the value
13def reg2(val, reg = 0.0):
@@
Figure 4 shows KL divergence from the initial policy and validation score at step 50 for ISOPO with p=1p=-1 (red) as described in section 2.0.1 vs sequence-wise euclidean normalization q=1q=-1 (yellow). While both normalizations increase the validation score, only the Fisher normalization has the effect of decreasing the KL drift.
Figure 4: KL-drift and validation score at step 50 for non-interacting ISOPO vs GRPO and REINFORCE (see section 4 for the setup). Multiple runs are shown for each algorithm and the convex hulls are shaded for visual clarity. ISOPO is shown with two different settings: with Fisher normalization p=1p=-1 described in section 2.0.1 (red) and with sequence-wise Euclidian normalization q=1q=-1 (yellow). While both increase the validation score, only the p=1p=-1 setting (red) decreases the KL drift.
@@ 3 Interacting variant
4 Results
I implemented ISOPO in VeRL [verl]. The experiments shown are for gsm8k with the Qwen-3 0.6B model. All runs use the group-relative advantage estimator, and all runs are without KL penalty. The two baselines shown are GRPO and REINFORCE. Here, GRPO refers to using PPO clipping of the importance sampling ratio, and REINFORCE refers to the same algorithm with no clipping. Like REINFORCE, ISOPO is run without PPO clipping. Figure 6 shows a comparison between best-of-3 validation curves for REINFORCE (green), GRPO (blue), and ISOPO (red) with q=1q=-1 and r=2r=-2. The code for ISOPO will be available at https://github.com/nilin/isopo.